### v4

# MIGE: Mutually Enhanced Multimodal Instruction-Based Image Generation and Editing

Abstract.
Despite significant progress in diffusion-based image generation, subject-driven generation and instruction-based editing remain challenging. Existing methods typically treat them separately, struggling with limited high-quality data and poor generalization. However, both tasks require capturing complex visual variations while maintaining consistency between inputs and outputs. Inspired by this, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions. It first treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation, then introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism. This unification enables joint training of both tasks, providing two key advantages: (1) Cross-Task Enhancement: by leveraging shared visual and semantic representations, joint training improves instruction adherence and visual consistency in both subject-driven generation and instruction-based editing. (2) Generalization: learning in a unified format facilitates cross-task knowledge transfer, enabling MIGE to generalize to novel compositional tasks, including instruction-based subject-driven editing. Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a SOTA in the new task of instruction-based subject-driven editing. Code and model have been publicly available at this link.
Controllable Image Generation, Diffusion Model, Joint Training
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755811isbn: 979-8-4007-2035-2/2025/10submissionid: 6167ccs: Computing methodologies Image representations
1. Introduction
Recent advances in diffusion models(Rombach et al., 2022; Peebles and Xie, 2023; Labs, 2023; Esser et al., 2024) have significantly advanced the development of customized image generation, enabling its widespread adoption in both subject-driven generation(Ruiz et al., 2023; Ye et al., 2023; Li et al., 2024a; Pan et al., 2023; Kang et al., 2025; Kim et al., 2024; Wu et al., 2025; Liang et al., 2025; Peng et al., 2024) and instruction-based editing(Brooks et al., 2023; Sheynin et al., 2024; Fu et al., 2023; Han et al., 2024). The former focuses on preserving a given subject, while the latter emphasizes flexible modifications guided by textual instructions. Their capability to generate personalized and controllable results demonstrates substantial practical value, thereby attracting growing research attention.
Existing approaches (Li et al., 2024a; Pan et al., 2023; Hu et al., 2024; Huang et al., 2024b) typically treat these two tasks independently and address them separately by tailoring specific input-output formats to each. However, this results in limited generalization capability and necessitates extensive training data. Specifically, subject-driven generation aims to preserve subject identity while adhering to textual instructions. Fine-tuning methods, like DreamBooth (Ruiz et al., 2023), require separate training for each subject during testing, which limits generalization. In contrast, zero-shot methods like KOSMOS-G (Pan et al., 2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs. Instruction-based editing, exemplified by InstructPix2Pix (Brooks et al., 2023), modifies images based on instructions. These methods require large, high-quality editing datasets to handle diverse instructions but struggle with maintaining local-to-global consistency.
The above separate processing methods can only rely on limited data and single-task optimization objectives, which hinders their ability to adapt to diverse subjects or editing requirements, and lack cross-task consistency, preventing a unified understanding of the tasks. To address these challenges, we shift our perspective from separate processing to the objectives of these two tasks. These two tasks share a fundamental principle: maintaining visual consistency between inputs and outputs while capturing variations guided by instructions. They also follow a similar input-output paradigm with complementarities: subject-driven generation emphasizes subject accuracy, while instruction-based editing manipulates images without altering irrelevant identities. These limitations and complementarities highlight the need for a unified framework that takes advantage of the strengths of both tasks to improve performance.
Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven image Generation and instruction-based Editing. The vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations. Structurally, we model subject-driven generation as creating an image on a blank canvas and instruction-based editing as modifying an existing image. This coherent input-output mapping simplifies the process while enabling both tasks to reinforce each other through joint training. Moreover, this integrated approach fosters emergent compositional capabilities beyond what either task can achieve alone.
Building on our unified framework, we address two critical challenges. First, for multimodal instruction encoding, existing methods (Li et al., 2023b; Pan et al., 2023; Li et al., 2024c) mainly rely on CLIP vision encoders (Radford et al., 2021) to extract semantic features, which is insufficient for preserving fine-grained subject details. To overcome this limitation, we propose a multimodal encoder with a feature fusion mechanism that integrates VAE (Kingma, 2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics. Secondly, we enhance compositionality through joint training, facilitating instruction-based subject-driven editing. To further optimize performance in this complex task, we develop a novel data construction pipeline based on a Multimodal Large Language Model (MLLM). This pipeline autonomously generates diverse multimodal instructions and corresponding output images. Moreover, to address the absence of an evaluation benchmark for this new composition task, we introduce MIGEBench. This specialized benchmark evaluates compositionality in terms of subject preservation and instruction adherence, providing a comprehensive assessment of the new task.
Our experiments demonstrate that joint training under the MIGE framework enables mutual enhancement between subject-driven generation and instruction-based editing, leading to significant performance gains over training on individual tasks. As shown in Figure 2, joint training markedly improves input-output consistency across both tasks, especially in editing, with scores rising from 0.821 to 0.873. Moreover, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in emerging compositional tasks. Extensive case studies (Figures 78) further demonstrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness across diverse scenarios.
Figure 2. Joint training significantly improves input-output similarity across both tasks. We show the average DINO and CLIP-I scores for subject-driven generation and instruction-based editing under single-task and joint training.
In summary, our contributions include:
  • Promising perspective of task mutual enhancement: We shift our perspective from traditional separate processing to mutual enhancement. Then, we design a unified input-output format for joint training on subject-driven generation and instruction-based editing, validating that joint training enables task enhancement.
  • Novel framework for Compositional Capability: We propose MIGE, a unified framework integrating two mainstream tasks in controllable image generation. Joint training further enables the compositional capability of instruction-based subject-driven editing. We also introduce a novel data pipeline and MIGEBench for comprehensive evaluation.
  • Strong Performance with Extensive Experiments: Extensive results show that MIGE performs competitively on subject-driven generation and instruction-based editing, with scores rising from 0.821 to 0.873, and sets a new state-of-the-art in instruction-based subject-driven editing.
2. Related Work
In this section, we discuss the most relevant studies. A more comprehensive discussion can be found in the Supplementary Material.
Universal generative model
Unifying tasks in image generation remains challenging. Some methods (Li et al., 2025a; Huang et al., 2024a) focus on bridging image understanding and generation, such as SEED-X (Ge et al., 2024), Emu2 (Sun et al., 2024), and WeGen (Huang et al., 2025), which use autoregressive models with auxiliary diffusion modules. UnifiedMLLM (Li et al., 2024b) employs MLLMs to learn shared representations and route to task-specific models. However, these approaches either emphasize understanding over generation or rely on complex pipelines. Others aim to unify generation and editing. Pixwizard (Lin et al., 2024) uses text-only instructions and task vectors. Instruct-Imagen (Hu et al., 2024) supports multimodal inputs but performs poorly on editing. ACE (Han et al., 2024) is limited to editing via transformer-based conditioning. OmniGen (Xiao et al., 2024) cannot distinguish reference images. OmniControl (Tan et al., 2024) merges all conditions with noise, while UniReal (Chen et al., 2024c) adds finer control but requires complex prompts. RealGeneral (Lin et al., 2025) enhances attention, and OnePic (Tao et al., 2024) depends on multiple references and masks. Overall, prior work often sacrifices flexibility or generality, combining tasks without leveraging their intrinsic connections. In contrast, we propose MIGE, which jointly trains on subject-driven generation and instruction-based editing, revealing their mutual benefits.
Subject-Driven Image Editing
Subject-driven image editing enables modifications based on a user-specified subject, typically including addition or replacement. Existing methods often rely on multiple reference images or precise cues. DreamEdit (Li et al., 2023a) and DreamMix (Yang et al., 2024) require multiple images for fine-tuning and iterative inpainting. PBE (Yang et al., 2023) uses compressed reference embeddings, while TIGIC (Li et al., 2025c) embeds subjects into predefined backgrounds. MADD (He et al., 2024a) supports diverse controls via masks or coordinates. AnyDoor (Chen et al., 2024b), AnyEdit (Yu et al., 2024), MimicBrush (Chen et al., 2024a) and BlobCtrl (Li et al., 2025b) require masks or additional range blobs. FreeEdit (He et al., 2024b) is limited to instruction templates and class-aware features. Overall, these approaches depend on detailed inputs and lack flexibility in handling diverse instructions for controllable generation.
Figure 3. Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse scenarios.
Figure 4. Overall framework of MIGE. MIGE consists of two components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model for modeling input-output relationships. The encoder incorporates a feature fusion mechanism to integrate visual and semantic features from reference image.
3. Method
In this section, we first shift our view from separate training to mutual enhancement. Guided by this new perspective, we first introduce the unified framework MIGE, followed by details of the architecture design and joint training approach.
3.1. Unified framework MIGE
Current methods treat subject-driven generation and instruction-based image editing separately, hindering performance due to data limitations and poor generalization. Since both aim to preserve visual consistency while incorporating instructed changes, unifying them allows mutual improvement. Joint training on diverse data enhances subject preservation and instruction following beyond the capabilities of individual task-specific models.
Therefore, we propose MIGE to unify the two tasks for joint training. By using multimodal instructions as a unified task representation, it supports flexible combinations and provides multimodal guidance. Additionally, we use conditional input to structurally unify the tasks, enhancing visual consistency. Their combination not only provides rich visual and instructional information but also naturally represents different tasks. We demonstrate the input-output format of MIGE in Figure 3.
3.1.1. Unified Multimodal Instruction
To enable joint training across multiple tasks, a unified task representation is essential. We introduce multimodal instructions composed of interleaved images and text, which provide both visual references and textual guidance for various controllable generation tasks. As shown in Figure 3, “¡imagehere¿” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions. This unified approach effectively accommodates subject-driven generation and instruction-based editing while remaining extensible to the more complex combination task.
3.1.2. Unified Conditional Input
We design a unified conditional input format to distinguish tasks while enabling shared capabilities. As shown in Figure 3, conditional inputs vary across tasks. For instruction-based editing, the VAE-encoded source image is used to preserve fine details of the original image, inform the model to edit only the target region, as done in prior editing methods (Brooks et al., 2023; Zhang et al., 2024). To standardize the input format, we concatenate an all-zero tensor for subject-driven generation to indicate generation from scratch with visual constraints from the multimodal instruction. This design unifies the input format, ensures input-output consistency, and facilitates extensibility to new tasks within a unified framework.
3.2. Architecture
The architecture of MIGE, as shown in Figure 4, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model (Peebles and Xie, 2023) for modeling input-output relationships. The diffusion model takes concatenated latent noise and conditional input along the channel dimension as input, performing controllable generation under the control of multimodal conditions. To further enhance the integration of visual and semantic information from the reference image, we introduce a novel feature fusion mechanism in the encoder.
3.2.1. Multimodal Encoder
To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder (Kingma, 2013) for visual feature extraction, a pretrained ViT from EVA-CLIP (Fang et al., 2023) for semantic feature extraction, a Q-Former (Li et al., 2023b), and a linear projection layer. Each image is represented by 32 tokens, which are then encoded by the LLM alongside text tokens to form a unified multimodal condition.
Prior works (Li et al., 2024a; Pan et al., 2023; Li et al., 2024c) focus on extracting semantics from reference images but lack the fine-grained details needed to preserve subject-specific features. To address this, we propose a feature fusion mechanism that combines the strengths of different visual encoders. ViT serves as a semantic feature extractor, while a VAE encoder acts as a visual feature extractor, leveraging its ability to compress and reconstruct images. As shown in Figure 4, we use the ViT semantic features compressed by the Q-Former, denoted as $f_{s}$, as guidance to adaptively incorporate the visual features $f_{v}$ extracted by the VAE. Inspired by TokenPacker (Li et al., 2024d), we represent rich features with fewer tokens, and apply ReLU (Glorot et al., 2011) to enhance stability. The fusion mechanism is expressed as: $f_{\text{img}}=f_{s}+\text{MLP}(\text{Attn}(f_{s},f_{v}))$ where $\text{Attn}(f_{s},f_{v})=\frac{Q(f_{s})\cdot K(f_{v})}{\sqrt{d}}V(f_{v})$. Here $d$ denotes the output dimension of the features, $Q$ is a linear layer, and both $K$ and $V$ are two-layer networks, where $K(\cdot)=V(\cdot)=\text{Linear}(\text{GELU}(\text{Linear}(\cdot)))$.
This fusion mechanism captures both semantic and visual information without introducing extra image tokens. The multimodal encoder compresses each image into 32 fused tokens, which replace the “¡imagehere¿” placeholder in the interleaved instruction. The LLM then encodes the entire prompt into a unified vision-language representation to guide the diffusion model.
(a) Subject Addition Dataset Construction Pipeline
(b) Subject Replacement Dataset Construction Pipeline
Figure 5. Data construction pipelines for instruction-based subject-driven image editing.
3.3. Joint Training
Multimodal instructions and conditional input unify task representation and input-output formats, enabling joint training. We fine-tune MIGE on data from all tasks to enhance cross-task synergy. Except for the two image encoders, all parameters are jointly trained to align the conditional space of the diffusion model with the multimodal encoder, as shown in Figure 4. This approach improves task coordination and consistency across modalities.
3.3.1. Data Construction
Joint training enables multi-task learning, balancing subject preservation and instruction, followed by modeling task relationships. We create a multi-task dataset for joint multimodal instruction tuning, covering subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image generation.
For subject-driven image generation, we follow the data construction methods of KOSMOS-G (Pan et al., 2023) and UNIMO-G (Li et al., 2024c), using an LLM to extract entities from captions, which are then fed into Grounded SAM (Ren et al., 2024) for segmentation. Subjects200k dataset from OmniControl (Tan et al., 2024) is also incorporated for better object preservation. For instruction-based editing, we filter existing datasets and use a rule-based strategy to construct multimodal instructions.
Instruction-based subject-driven image generation is an emerging task involving two subtasks: instruction-based subject addition and subject replacement, which allows users to add or replace a specified subject in an image using multimodal instructions. However, there is no sufficient dataset available for this task.
For instruction-based subject addition, we propose a pipeline inspired by SAM-FB (He et al., 2024a), as shown in Figure 5(a). Starting with the SA-1B (Kirillov et al., 2023) dataset, we construct the source image and a multimodal instruction for input-output pairs. We use SAM (Kirillov et al., 2023) to segment annotated entities, filter and retain the main subject using an MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o (Hurst et al., 2024). Due to resource limitations, we process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more. For instruction-based subject replacement, we filter from existing editing data, use Grounded SAM for subject segmentation, and construct multimodal instructions to form input-output pairs, as shown in Figure 5(b). We also introduce virtual try-on data constructed using IDM-VTON (Choi et al., 2024), resulting in approximately 110k samples. More details on training data construction are in Supplementary Material.
4. Experiments
In this section, we present experimental results on three tasks, followed by analysis of the joint training strategy and framework design. For subject-driven image generation, we compare our method with both task-specific models using MLLMs as encoders and recent unified frameworks. For instruction-based image editing, we similarly compare against both task-specific and unified models, including SmartEdit (Huang et al., 2024b) and ml-MGIE (Fu et al., 2023), which also use MLLMs. For instruction-based subject-driven image editing, given the lack of instruction-only baselines, we additionally evaluate mask-dependent methods using masks from MIGEBench.
4.1. Implementation Details
[…]
4.2. Evaluation Results
As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models. MIGE achieves strong performance in subject-driven generation and instruction-based editing, and shows promising results on instruction-based subject-driven image generation. Detailed results are in Supplementary Material.
Table 1. Quantitative results on the Emu Edit and MagicBrush test sets. MIGE outperforms other universal models on MagicBrush, demonstrating superior instruction alignment and detail preservation.
Emu Edit Test set MagicBrush Test set
[…]
Task-specific
Models
InstructPix2Pix (Brooks et al., 2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
MagicBrush (Zhang et al., 2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
UltraEdit (Zhao et al., 2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
SmartEdit (Huang et al., 2024b) 0.752 0.830 0.290 0.100 0.129 0.055 0.821 0.879 0.301 0.122 0.084 0.033
ml-MGIE (Fu et al., 2023) 0.779 0.843 0.276 0.040 0.101 0.034 0.751 0.844 0.274 0.039 0.102 0.034
Unified
Models
PixWizard (Lin et al., 2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
ACE (Han et al., 2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
OmniGen (Xiao et al., 2024) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
OneDiff (Le et al., 2024) 0.728 0.819 0.285 0.082 0.136 0.060 0.742 0.840 0.295 0.107 0.128 0.052
MIGE (ours) 0.832 0.865 0.306 0.114 0.088 0.027 0.889 0.905 0.306 0.119 0.055 0.013
/only_edit data 0.785 0.841 0.302 0.104 0.117 0.046 0.796 0.862 0.300 0.111 0.094 0.036
/wo_VAE feature 0.799 0.846 0.300 0.098 0.103 0.035 0.811 0.868 0.299 0.105 0.098 0.036
/wo_multimodal instruction 0.592 0.723 0.280 0.113 0.177 0.074 0.548 0.714 0.277 0.135 0.170 0.066
4.2.1. Instruction-based Image Editing
Instruction-based image editing allows users to modify a source image using free-form multimodal instructions, such as adding, removing, or altering objects and styles. Table 1 reports results on Emu Edit and MagicBrush (Zhang et al., 2024). DINO and CLIP-I assess similarity to the source image, CLIP-T evaluates alignment with the target caption, CLIPdir measures consistency between text and image embedding changes, and L1/L2 capture pixel-level differences.
As shown in Table 1, MIGE achieves the highest CLIP-T and outperforms all task-specific models on CLIPdir, indicating stronger instruction following. On MagicBrush, it surpasses all universal models across almost all metrics, with the best DINO, CLIP-I, CLIP-T, and lowest L1/L2, demonstrating superior instruction fidelity and detail preservation. This capability is further illustrated in Figure 6, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas. Compared to ml-MGIE, which also uses a MLLM as an encoder, MIGE better preserves input-output consistency and avoids altering irrelevant backgrounds.
Figure 6. Qualitative comparison for instruction-based image editing. MIGE demonstrates superior editing accuracy and instruction understanding.
4.2.2. Subject-driven Image Generation
Generating images that satisfy both image and text constraints from a multimodal prompt is challenging. We compare MIGE with three task-specific methods using MLLMs as encoders and three universal models on DreamBench (Ruiz et al., 2023). DINO and CLIP-I measure subject fidelity, while CLIP-T assesses instruction adherence. As shown in Table 2, MIGE surpasses all unified models in subject preservation ability, as evidenced by the highest DINO and CLIP-I scores, with the DINO score even outperforming task-specific models. Examples in Figure 7 further highlight MIGE’s ability to effectively preserve the identity of each object in multi-object scenarios and combine them based on instructions, demonstrating that our multimodal encoder effectively captures both semantic and visual features simultaneously.
Table 2. Quantitative results for subject-driven image generation on DreamBench. Results with * denote our reproduction. MIGE outperforms universal models in subject preservation and is competitive with task-specific models.
Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Task-specific
Models
BLIP-Diffusion (Li et al., 2024a) 0.670 0.805 0.302
KOSMOS-G (Pan et al., 2023) 0.694 0.847 0.287
UNIMO-G (Li et al., 2024c) 0.668 0.841 0.329
Unified
Models
Omnigen * (Xiao et al., 2024) 0.711 0.800 0.312
UniReal (Chen et al., 2024c) 0.702 0.806 0.326
OneDiffusion * (Le et al., 2024) 0.582 0.750 0.240
MIGE (ours) 0.744 0.830 0.293
/only_subject data 0.726 0.823 0.289
/wo_VAE feature 0.741 0.828 0.287
Figure 7. Qualitative comparison for subject-driven image generation. MIGE exhibits superior subject preservation, even in multi-subject scenarios.
Table 3. Quantitative results on instruction-based subject-driven editing. Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions. Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
(a) Results on instruction-based subject replacement.
Methods Editing Subject Preserving
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
I-O sim 0.668 0.842 0.271
Mask-denpendent
Models
PBE (Yang et al., 2023) 0.810 0.885 0.304 0.521 0.792
AnyDoor (Chen et al., 2024b) 0.817 0.890 0.307 0.551 0.789
TIGIC (Li et al., 2025c) 0.789 0.874 0.313 0.453 0.744
MADD (He et al., 2024a) 0.736 0.852 0.284 0.446 0.742
Mask-free
Models
OmniGen (Xiao et al., 2024) 0.802 0.868 0.296 0.580 0.792
MIGE(ours) 0.863 0.909 0.307 0.652 0.834
/subject data + edit data 0.789 0.873 0.299 0.503 0.764
/only_compositional data 0.780 0.860 0.303 0.585 0.808
/wo_VAE feature 0.852 0.900 0.308 0.633 0.819
(b) Results on instruction-based subject addition.
Methods Editing Subject Preserving
DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
I-O sim 0.783 0.903 0.295
Mask-denpendent
Models
PBE (Yang et al., 2023) 0.843 0.908 0.321 0.495 0.794
AnyDoor (Chen et al., 2024b) 0.863 0.930 0.324 0.533 0.799
TIGIC (Li et al., 2025c) 0.840 0.901 0.325 0.455 0.753
MADD (He et al., 2024a) 0.885 0.930 0.316 0.519 0.785
Mask-free
Models
OmniGen (Xiao et al., 2024) 0.791 0.870 0.312 0.605 0.814
MIGE(ours) 0.909 0.940 0.322 0.638 0.838
/subject data + edit data 0.807 0.895 0.304 0.309 0.683
/only_compositional data 0.879 0.928 0.324 0.577 0.820
/wo_VAE feature 0.902 0.934 0.322 0.633 0.831
4.2.3. Instruction-based Subject-driven Image Editing
Benchmark Construction
Instruction-based subject-driven image editing is a novel task. Existing methods rely on masks or positional cues for editing (Li et al., 2023a; Yang et al., 2024) but lack support for instruction-only editing. The AnyEdit-Test Benchmark (Yu et al., 2024) supports few instruction templates and evaluates only one type of object editing based on automatically extracted masks. Current benchmarks (Yang et al., 2023; Chen et al., 2024b) for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
To address these issues, we manual construct MIGEBench, a benchmark consisting of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement. More details can be found in the Supplementary Material.
Evaluation Results
Our evaluation focuses on editing ability and subject preservation. DINO and CLIP-I measure the similarity of the generated image to the ground truth, while CLIP-T evaluates alignment with the target caption, indicating editing ability. Subject preservation is evaluated by extracting the edited subject via Grounded SAM (Ren et al., 2024) and comparing it to the input subject image using DINO and CLIP-I. This separates the evaluation of image-level editing from subject-level feature preservation. Methods that do not support instruction-based editing use masks provided in MIGEBench during testing.
Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b). We compute the DINO, CLIP-I, and CLIP-T scores between the input image and the ground truth. The first row of the table (denoted as ”I-O sim”) provides a baseline for evaluating the effectiveness of editing ability. Mask-dependent models can easily preserve the edit-unrelated regions, but MIGE still outperforms all methods, highlighting the effectiveness of conditional inputs and MIGE’s precise understanding of instructions. The CLIP-T score measures instruction adherence in mask-free models. MIGE significantly outperforms OmniGen in both tasks, demonstrating the advantage of our unified approach. MIGE also achieves the best performance on subject preservation in both tasks. As shown in Figure 8, MIGE successfully understands the task of replacing a bird with an owl, rather than simply overlaying or adding an owl, demonstrating the effectiveness of the multimodal encoder which simultaneously encoding the semantic and visual information of the reference image.
Figure 8. Qualitative results on the benchmark for the subject addition and subject replacement. The upper section compares subject addition results, while the lower section compares subject replacement. During testing, the ¡imagehere¿ placeholder in the multimodal instruction is replaced according to the image sequence. MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
4.3. Further Analysis
4.3.1. Effectiveness of Joint Training
To assess the effectiveness of joint training, we train models separately on individual datasets (denoted as “only_subject data,” “only_edit data,” and “only_compositional data”) and compare their performance with the jointly trained model. The results in Tables 1, 2 and 3 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other. As shown in Table 3, joint training also enhances the performance of the compositional new task, further highlighting its overall benefits. These findings emphasize both the effectiveness and necessity of joint training. In conclusion, joint training of subject-driven generation and instruction-based editing within our unified framework not only boosts compositional capability but also improves the performance of each individual task. This validates our motivation and provides further theoretical support for unified models for multiple tasks.
4.3.2. Effectiveness of Feature Fusing
MIGE employs a feature fusion mechanism in the multimodal encoder to integrate semantic features from ViT and visual features from VAE. As shown in Table 2, Table 1 and Table 3, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing. This is particularly evident in the improved CLIP-I and DINO scores and the significant reduction in L1 and L2 metrics, demonstrating that the inclusion of additional visual features helps maintain consistency between the input and output.
4.3.3. Effectiveness of Instruction-based Subject-driven Image Editing Dataset
Joint training of subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”). To enhance the capability of MIGE in this new task, particularly in understanding spatial terms and size descriptions, we constructed a task-specific dataset for joint training. As shown in Table 3(a) and Table 3(b), the task-specific data significantly improved the model’s overall abilities. This demonstrates the effectiveness of our constructed dataset, and the proposed data generation pipeline serves as a valuable reference for future dataset construction.
4.3.4. Effectiveness of Multimodal Instruction
Existing instruction-based editing works (Brooks et al., 2023; Sheynin et al., 2024; Zhang et al., 2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions. To measure the benefit of multimodal instructions, we trained the model with text-only editing instructions for comparison. As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models. As shown in Table 1, using multimodal instructions consistently improves performance over text-only instructions (denoted as “wo_multimodal instruction”). This enhances input-output consistency and instruction-following ability in multi-task training. The significant improvement in the L1 and L2 metrics indicates finer control over images and more accurate edits. While text-only instructions provide the necessary changes, the high CLIPdir score and lower values in other metrics show that multimodal instructions add visual context, enabling more precise and faithful modifications. For more qualitative results, see Supplementary Material.
4.3.5. Comparison of Data Volume
Task-specific models treat tasks separately and typically require large amounts of high-quality data. In contrast, by exploiting the relationships between tasks and the unified structure, MIGE achieves excellent results with minimal data through joint training. As shown in Table 4, MIGE demonstrates that both the model structure and joint training are key to achieving this performance with minimal data, using fewer data than task-specific models and other unified models.
Table 4. Comparison of Data Volume Across Methods
  Methods   Data Volume
  Task-specific
  Models
  MADD (He et al., 2024a)   3M
  PBE (Yang et al., 2023)   16M
  KOSMOS-G (Pan et al., 2023)   3159.3M
  UNIMO-G (Li et al., 2024c)   304.3M
  Unified
  Models
  PixWizard (Lin et al., 2024)   30M
  UniReal (Chen et al., 2024c)   100M
  ACE (Han et al., 2024)   700M
  Omnigen (Xiao et al., 2024)   326.06M
  MIGE (ours)   2.28M
5. Conclusion
We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing, leveraging multimodal instructions and conditional input for joint training that enhances task synergy. Joint training also unlocks new capabilities like instruction-based subject-driven image editing. We introduce pipelines for this new task to construct training data and MIGEBench for evaluation. Our experiments show that joint training leads to significant enhancement in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks. This integration enhances controllability and offers promising directions for future multimodal image generation and editing.
6. Supplementary Material
Appendix A Qualitative Comparison
We conduct a qualitative comparison to assess the visual performance and effectiveness of MIGE. Figure 14 presents a range of cases on each task to showcase the generation results of MIGE, demonstrating its ability to handle diverse inputs and produce high-quality outputs. This comparison highlights the strengths of MIGE and the benefit of joint training, offering valuable insights into its potential for real-world applications.
Appendix B Related Work
B.1. Subject-Driven Image Generation
Subject-driven generation focuses on preserving given subject features while generating new images. DreamBooth (Ruiz et al., 2023) fine-tunes the full model, leading to high resource costs. IP-Adapter (Ye et al., 2023) trains lightweight attention layers, struggles with prompt-image balance. Methods like Blip-Diffusion (Li et al., 2024a) use embeddings to condition generation, yet face challenges in subject preservation and generalization. KOSMOS-G (Pan et al., 2023) and UNIMO-G (Li et al., 2024c) leverage Multimodal Large Language Models (MLLMs) as flexible encoders, but require large datasets for task-specific adaptation and lack robust instruction-following ability, limiting transferability.
B.2. Instruction-based Image Editing
Instruction-based editing modifies images based on instructions, offering greater flexibility than caption-driven methods. Inst-inpaint (Yildirim et al., 2023) and EraseDraw (Canberk et al., 2025) focus on object-level edits. InstructPix2Pix (Brooks et al., 2023), EmuEdit (Sheynin et al., 2024), and MagicBrush (Zhang et al., 2024) enable diverse edits via large-scale data. AnyEdit (Yu et al., 2024), FireEdit (Zhou et al., 2025), and TF-TI2I (Hsiao et al., 2025) introduce task-aware routing, region prompts, or fixed templates. Recent works like InsightEdit (Xu et al., 2024), MGIE (Fu et al., 2023), and SmartEdit (Huang et al., 2024b) leverage MLLMs to enhance instruction understanding. However, maintaining input-output consistency remains challenging.
Appendix C Dataset Construction
To enable joint training of MIGE, we designed a series of pipelines for data construction and processing. Our dataset includes three tasks: subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image editing. The proportions of the training data are shown in Figure 9. In this section, we will detail the pipelines for processing each data component and other related details.
Figure 9. Composition of training data.
C.1. Subject-driven Image Generation Data Construction
We construct multimodal instruction-image pairs by replacing entities in image captions with corresponding entity images. The dataset combines BLIP3-GROUNDING-50M (Xue et al., 2024) and internal sources. For the former, we filter out low-quality pairs with CLIP-T scores below 0.255 and retain entities with bounding boxes sized between 0.05 and 0.8. Entities are segmented using SAM (Kirillov et al., 2023), and pairs with low CLIP-T scores between the prompt ”a photo of [entity]” and the segmented image are removed. For internal data, we adopt the KOSMOS-G’s pipeline, using MPT-7B-Instruct (Team et al., 2023) to extract entities. Entity images are obtained via Grounded SAM (Ren et al., 2024), and filtered by a CLIP-T score threshold of 0.255 and a size ratio between 0.1 and 0.85. To enhance subject fidelity, we further include 112,846 samples from subject200k, following OminiControl (Tan et al., 2024).
C.2. Instruction-based Image Editing Data Construction
Previous works on instruction-based editing relied solely on text, covering object-level additions, deletions, modifications, and global changes in background or style. We integrated data from InstructPix2Pix (Brooks et al., 2023), UltraEdit (Zhao et al., 2024), MagicBrush (Zhang et al., 2024), SEED-Data-Edit (Ge et al., 2024), GIER (Shi et al., 2020), and HQ-Edit (Hui et al., 2024), filtering out instances with target image aesthetic scores above 5.5. To enhance multimodal instruction alignment, we reformulated text-based instructions based on task type. Specifically, for instructions containing phrases like “in the image” or “to the image,” we replaced them with “in¡imagehere¿” or “to¡imagehere¿.” For others, we appended “of¡imagehere¿” for global edits or “in¡imagehere¿” for localized modifications.
C.3. Instruction-based Subject-driven Image Editing Data Construction
For subject addition, inspired by SAM-FB (He et al., 2024a), we construct background images, foreground entities, and multimodal instructions, as shown in Figure 5(a). We filter SA-1B (Kirillov et al., 2023) images with aesthetic scores above 5 and select entities with a foreground-to-image ratio between 0.1 and 0.5. Qwen2-VL-7B (Wang et al., 2024) verifies entity completeness. The selected entities serve as foregrounds, and backgrounds are inpainted using LAMA (Suvorov et al., 2022). Foreground-related text is extracted with Qwen2-VL-7B, and multimodal instructions are generated via GPT-4o (Hurst et al., 2024). Due to resource constraints, we process the first 500,000 SA-1B samples, yielding 193,247 multimodal instruction-foreground-background pairs.
For subject replacement, we construct foreground entities and multimodal instructions, as shown in Figure 5(b). SEED-Data-Edit’s part1-unsplash is filtered to retain replace-task samples with annotation match scores above 0.3 and aesthetic scores above 5.5. Qwen2.5-14B-Instruct (Team, 2024) verifies the main entity. The target entity is replaced with <imagehere> to form the multimodal prompt. Grounded SAM segments entities based on annotations, resulting in 79,693 multimodal instruction-replace entity-source image-target image pairs. We also include 34,947 virtual try-on samples generated using IDM-VTON (Choi et al., 2024).
Appendix D Prompts
LLMs and MLLMs are essential to our data and benchmark construction, with all prompts listed in Table 6 to Table 9.
Appendix E Implementation Details
E.1. Training Details
[…]
MIGE is trained on our multi-task dataset using the AdamW optimizer (Loshchilov, 2017) with a weight decay of 0.03 and a learning rate of 1e-5 for 18 epochs on 48 H20 GPUs, totaling six days of training with a batch size of 960 (20 per GPU).
E.2. Evaluation Details
We evaluated the subject-driven generation capability of MIGE on DreamBench, which contains 750 prompts covering 30 subjects. For each prompt, four images were generated using seeds 0, 1, 2, and 3, resulting in 3,000 images. Following KOSMOS-G, we selected one image per prompt. Then we use the extracted subject from Grounded SAM as input, aligning with the training process. Subject fidelity was assessed using DINO and CLIP-I, while CLIP-T measured adherence to multimodal instructions.
For instruction-based editing evaluation, the Emu Edit benchmark contains known issues, including incorrect image-caption pairs and duplicate source-target captions. Prior works have handled these inconsistencies differently, leading to incomparable results. Therefore, we reimplemented all currently available open-source methods on both test sets using a fixed random seed (seed=0) for consistency. Consequently, CLIP-I, L1, and DINO were computed on 3,589 Emu Edit test samples, while CLIPdir and CLIP-T were evaluated on 2,899 samples after removing problematic data.
Figure 10. The pipeline of benchmark construction.
Appendix F Benchmark Construction
Instruction-based subject-driven image editing is a novel task lacking evaluation benchmarks. Previous works (Li et al., 2023a; Yang et al., 2024) did not support variety of instructions and lacked ground truth target images. To address this, we designed a benchmark construction pipeline with multiple manual inspections to ensure image and caption quality.
As shown in Figure 10, we first filter valid editing pairs using a Gradio (Abid et al., 2019) interface, as shown in Figure 11. Then, Qwen2.5-14B-Instruct extracts entity names and constructs multimodal instructions. After manually reviewing object names, we use Grounding DINO (Liu et al., 2025) for bounding boxes and SAM for segmentation. Cropped entities are saved with black and white backgrounds for diverse test scenarios. Prompts are detailed in Supplementary Material D.
For evaluation, Qwen2-VL-7B generates captions comparing generated and ground truth images, followed by manual reviews to remove unrecognized entities and correct inaccuracies. The final benchmark includes 500 samples each for subject addition and replacement. By avoiding artifacts from inpainting and incorporating multiple manual reviews, we ensure high-quality data. Figure 12 and Figure 13 showcase some example benchmark pairs.
Figure 11. The annotation interface for benchmarks built with Gradio.
Figure 12. Subject addition examples in our benchmark.
Figure 13. Subject replacement examples in our benchmark.
Our benchmark, MIGEBench, is constructed based on SEED-Data-Edit, which is diversity. The benchmark covers a total of 565 distinct subjects. For the subject-replacement subset, the data is distributed across 10 major, well-balanced categories, as shown in 5. This hierarchical and balanced structure ensures a high degree of diversity throughout the MIGEBench dataset.
Table 5. Category-wise Statistics of MIGEBench
Category Subject Replacement Subject Addition
Animals 62 77
People 60 66
Vehicles 40 52
Buildings & Structures 27 50
Plants & Nature 30 33
Household & Man-made Objects 88 91
Art & Decor 54 47
Food & Drink 23 8
Clothing & Accessories 31 24
Other 85 52
Total 500 500
Table 6. Prompt design for subject completeness evaluation in an image. The model should identify whether the subject is complete, not abstract elements.
Prompt for Qwen2-VL-7B to determine the completeness of subject in the image
Prompt: Determine if the subject in the image is complete. If it is complete and not an abstract object such as background, grass, sky, tree, stone, or part of another item, please return True. Otherwise, return False.
Table 7. Prompt design for GPT-4o to generate multimodal instruction for subject addition training data.
Prompt for GPT-4o to generate multimodal instruction
Prompt: The object is {object_name}. It is located in a bounding box with coordinates ({x}, {y}, {w}, {h}) on an image of size {width}x{height}. Describe its size, relative position, and relation to surrounding objects. Avoid describing the overall scene or unrelated elements.
Your response should start with “Add <imagehere> to the [position] of <imagehere>”. (The first <imagehere> indicates the object and the second indicates the image.) Keep the <imagehere> symbol in the first sentence in your reply.
Answer briefly in two sentences:
Table 8. Prompt design for determining if an edit instruction pertains to the main subject of the image. The model evaluates whether the instruction relates to humans, animals, or other environmental features.
Prompt for Qwen2.5-14B-Instruct to determine whether an edit instruction pertains to the main subject of an image
Prompt: You are an assistant that determines whether an edit instruction pertains to the main subject of an image.
The main subject refers to humans or animals only. If the edit instruction is related to the main subject, respond with ’yes’.
If it pertains to background, large areas of vegetation, or environmental information, respond with ’no’.
Here are some examples:
1. replace the grass with sand $\rightarrow$ no
2. replace the trees with palm trees $\rightarrow$ no
3. replace the dirt road with a cobblestone path $\rightarrow$ no
4. replace the bird with a parrot $\rightarrow$ yes
Now, analyze the following instruction and respond accordingly: {instruction}
Table 9. Prompt designs for extracting and modifying edit instructions based on either replacement or addition of subjects.
Prompt for extracting the main subject and modifying the edit instruction
Prompt 1: Replace Examples
You are an assistant that determines the main subject of an edit instruction and outputs the extracted main subject along with a modified prompt.
Given an edit instruction like “replace X with Y”:
- Extract “Y” (the replacement subject).
- Output the extracted subject.
- Modify the instruction by replacing “Y” with <imagehere> in the original prompt and output the new instruction.
Examples:
1. replace one woman with a man $\rightarrow$ Output: a man, replace one woman with <imagehere>
2. replace the castle with another castle $\rightarrow$ Output: castle, replace the castle with <imagehere>
3. replace the desk with a white one in the bottom middle $\rightarrow$ Output: desk, replace the desk with <imagehere> in the bottom middle
Prompt 2: Add Examples
You are an assistant that extracts the added subject in the edit instruction.
Given an edit instruction like “add X to the image”:
- Extract “X” (the added subject).
- Output the extracted subject.
- Modify the instruction by replacing “X” with <imagehere> in the original prompt and output the new instruction.
Here are some examples:
1. add human over the stone in the bottom left $\rightarrow$ Output: human, add <imagehere> over the stone in the bottom left
2. add a car on the road at the bottom $\rightarrow$ Output: a car, add <imagehere> on the road at the bottom
3. Add an owl on the left shoulder $\rightarrow$ Output: an owl, Add <imagehere> on the left shoulder
Figure 14. Qualitative results of subject-driven image generation (top) , instruction-based image editing (middle), and instruction-based subject-driven image editing (bottom).
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No.62172393), and Major Public Welfare Project of Henan Province (No.201300311200).